A particle of mass m is executing uniform circular motion on a path radius r if P is the magnitude of its linear momentum The radial force acting on the particle is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

given P =mv

 so v=P/m

so radial force is =  mv2/r

                            =  m P2 /m2 r

                             = P2/mr

 Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin